OpenStudy (anonymous):

y=-9x2,y=1/3x2,y=1/2x2

OpenStudy (imstuck):

what are you doing with these?

OpenStudy (anonymous):

\[{y = -9×2, y = \frac{ 1 }{ 3}×2, y = \frac{ 1 }{ 2 }×2}\]

OpenStudy (anonymous):

2